The electronic configuration of an atom's valence shell is ns2np2. During the titration of a CuCl2 with conc. H2SO4(0.1M) acid and a suitable indicator was used such that PH(y) of the HCl was 1.500. What is the period to which the element belongs with respect to the fact that y= -log(10+10n/2)



1
Expert's answer
2021-11-11T00:41:02-0500

Y= log (10+101.5/2)

Y= log (10+100.75)

Y= log (10+5.62)

Y= log 15.62= 1.19

Hence belongs to period 1